# RING-NeRF: A Versatile Architecture based on Residual Implicit Neural Grids

### v1

@@ Abstract
Since their introduction, Neural Fields have become very popular for 3D reconstruction and new view synthesis. Recent researches focused on accelerating the process, as well as improving the robustness to variation of the observation distance and limited number of supervised viewpoints. However, those approaches often led to dedicated solutions that cannot be easily combined. To tackle this issue, we introduce a new simple but efficient architecture named RING-NeRF [^1] The source code of RING-NERF will be released on Github., based on Residual Implicit Neural Grids, that provides a control on the level of detail of the mapping function between the scene and the latent spaces. Associated with a distance-aware forward mapping mechanism and a continuous coarse-to-fine reconstruction process, our versatile architecture demonstrates both fast training and state-of-the-art performances in terms of: $(1)$ anti-aliased rendering, $(2)$ reconstruction quality from few supervised viewpoints, and $(3)$ robustness in the absence of appropriate scene-specific initialization for SDF-based NeRFs. We also demonstrate that our architecture can dynamically add grids to increase the details of the reconstruction, opening the way to adaptive reconstruction.
Figure 1: Performance comparisons for different tasks of our architecture RING-NeRF against the corresponding SotA solutions: Nerfacto [18] / NeuS-facto [20] for fast reconstruction, ZipNeRF [3] for anti-aliased novel view synthesis, FreeNeRF [19] for reconstruction from few supervision views (9 views are used here), and NeuralAngelo [10] for robustness of SDF reconstructions without scene-specific initialization. RING-NeRF achieves on-par performances with dedicated solutions with much faster reconstruction speed. PSNR values (masked object for the few images setup and estimated on the RGB renders for the SDF) for each images and training reconstruction times are inset.
@@ 1 Introduction
Neural Radiance Fields (NeRF) have emerged as a novel method for representing 3D scenes using neural networks. In its original design [12], a simple multi-layer perceptron (MLP) is trained to reproduce a continuous 5D function that outputs the density and radiance emitted in each direction ($\theta$, $\phi$) at each point (x, y, z) in space. This approach inspired many works due to the impressive quality of its novel view synthesis as well as its simplicity.
However, the so-called ”vanilla” NeRF architectures converge very slowly as they use large neural networks. Instant-NGP (I-NGP) [13] introduced a new architecture based on a hierarchical hash-grid pyramid of learnable feature codes describing the 3D scene and combined with a much shallower MLP, called decoder, that transforms the concatenation of the codes, interpolated from the grids, into density and radiance. Resulting in more local updates, this method reduced the training duration from hours to minutes and swiftly became the second standard architecture for Neural Fields.
@@
Nature of scene - by transitioning from object-centric scenes to open unbounded scenes, using, for instance, space contraction [2];
Robustness - by managing free motion trajectories with variations of the observation distance while avoiding aliasing artifacts, mostly through the integration of Level of Detail (LOD) in the model [1, 3], or by reducing drastically the number of supervised views through different kinds of regularization [19, 7, 14];
Extensibility - by going from a holistic reconstruction to an incremental process, mainly through the use of a frozen decoder [23].
However, most of these solutions introduce specific and complex mechanisms that both increase the computational cost and the difficulty of fusing them together.
In this article, we demonstrate that representing a 3D scene as a multi-scale mapping function from 3D scene space to a decoder latent space brought us to define RING-NeRF, a simple grid-based architecture that improves several areas of NeRF without the need for a specific solution for each issue. It provides state-of-the-art-performances in terms of quality, speed, robustness and extensibility while keeping the overall process simple to implement and use. Our contributions can be summarized as follows:
@@
    an architecture that, by construction, represents the scene with a continuous level of detail, and enables spatial and resolution extensibility;
@@
    a distance-aware forward mapping that avoids aliasing artifacts while not slowing down the reconstruction;
@@
    a continuous coarse-to-fine reconstruction process that improves the convergence and stability in challenging setups (supervision with few viewpoints, no scene-specific initialization for SDF reconstruction).
2 Related Work and Motivation
@@
From its original version [12], the great majority of current researches focused on overtaking limitations of NeRF-based reconstruction in terms of adaptability to various scene natures, robustness to observation distance or to limited amount of viewpoints and in terms of reconstruction extensibility.
@@
Adaptability to various natures of the scene. The ability of Neural Fields to reconstruct various natures of scenes depends on different factors. The first one is related to its architecture itself. Tri-plane architectures [6, 24] are mostly designed for object-centric reconstruction while ”vanilla” and 3D Grid-based NeRF are able to cover a wider variety of scene even though they initially were still constrained to a limited volume. A second factor is related to the representation of the 3D space of the scene, especially to represent the background in open scenes. While some approaches use two different NeRF models, to reconstruct separately the foreground and the background [22], others apply space contraction to the 3D scene coordinates [22, 2, 3] to map the infinite scene volume into a bounded one. The last factor is related to the initialization of the Neural Fields. While initialization of density-based NeRF can adjust to almost any nature of scenes, the convergence of SDF-based (Signed Distance Function) Neural Fields is extremely sensitive to their parameters initialization. For the latter, current solutions rely on a scene-specific initialization (using a SDF field representing a sphere for object-centric scenes or an inverted sphere for open scenes), making the solution unable to adapt automatically to any scene.
@@
Robustness to observation distance variations. The initial NeRF model, as well as most of the subsequent works, relies implicitly on the hypothesis of a constant distance of observation to the scene. Indeed, while based on volume rendering, the NeRF model provides a per-3D-point scene density representation and a rendering process which doesn’t take into account the increasing volume covered by a pixel with respect to the distance to the cameras. As underlined in Mip-NeRF [1], this discrepancy induces artifacts, such as over-contrasted images or aliasing phenomenon, when the distance of observation or image resolution differs from the ones used at reconstruction time. To avoid these artifacts, the rendering process needs to assess the density and color for a volume instead of a point. Replacing the per-point density/color with a per-volume implies to represent the scene with different levels of detail (LOD), in order to vary the precision of the reconstruction based on the observation distance. A first solution consists of using a LOD-aware latent space, meaning that the LOD information is already encoded in the inputs of the MLP. This is usually achieved by using a per-LOD decoder [4, 16], or by downweighting the components of the latent feature [1, 3], or even by augmenting the feature with LOD information [3]. A second approach consists of defining the latent representation of a volume as the mean of the latent features of the points included in the volume. While keeping the latent space unaware of the LOD, it requires integrating the features over the volume. Such computation can be achieved through convolutions [6, 24] for tri-plan representation, or through super-sampling of the latent space for 3D grids [3].
@@
Robustness to limited amount of viewpoints. By construction, NeRF is subject to the shape-radiance ambiguity [22]. If not enough supervision viewpoints are available, the optimization might overfit those viewpoints while not providing consistent 3D reconstruction nor generalization to non-supervised viewpoints. A first family of solutions to overcome this limitation consists of regularizing the reconstruction process through additional losses, whether via geometric [14, 21] or semantic [7] regularization. The second family of solutions relies on a progressive reconstruction of the details of the scene [19]. Restraining the ability of the model to reconstruct a complex scene at early stages enforces the consistency of the reconstruction over the different supervision viewpoints. This presents the advantage of bringing stability while keeping a fairly simple training process. It is to be noted that all of these methods rely on vanilla Neural Fields and to the best of our knowledge, no grid-based methods have been introduced yet.
@@
Reconstruction Extensibility. Two kinds of extensibility should be distinguished: extensibility with respect to number of views or to the resolution. The first one is related to reconstructing new scene areas while keeping the previously reconstructed ones unchanged. When the number of supervision views becomes large, this problem becomes intractable for vanilla NeRF architecture since any update of the MLP affects the whole reconstruction. The usual solution consists of using grid-based approaches with a pre-trained decoder that is frozen during the reconstruction[23]. On the other side, extensibility of the resolution consists in continually refining the level of detail of a previously reconstructed scene. This problem is intractable for classic architectures such as vanilla NeRF and I-NGP since the number of layers (resp. grids) of those approaches cannot increase during the reconstruction. Some rare solutions, such as Neural Sparse Voxel Fields [11], combine a pre-trained decoder with a data structure allowing to dynamically allocate voxels to increase the reconstruction resolution.
Motivation. While each of these previous works provided significant improvement for a specific use case, their architecture and process pipeline were designed to best perform on this specific application without considering other ones. Based on this previous knowledge, we propose RING-NeRF, a simple architecture and reconstruction process that is designed to improve its versatility. It relies on:
  • Continuous and non-destructive Level of Detail: LOD-aware representation provides robustness to observation distance variation and robustness to limited supervision views through a coarse-to-fine reconstruction process. Combining both advantages, adding finer LOD during the reconstruction should not damage the coarser ones, unlike [4, 10];
  • Position and LOD invariance of latent space: extensibility relies on the ability of a pre-train decoder to generalize over new scenes. To favor this generalizability of the decoder latent space, the latter should not possess any information relative to the position of the sample (unlike [16, 24]) nor its associated LOD, unlike [1, 2, 3, 16, 4, 10].
@@ 3.1 Overview
Figure 2: Overview of RING-NeRF: to render a pixel, the casted cone is sampled with cubes. Depending on the cube volume, the corresponding LOD of the scene is selected and the latent feature is computed from the Residual Implicit Neural Grids accordingly. The density and color of the cube are first decoded from the latent feature with a tiny MLP and then integrated with other samples through volume rendering.
Our objective is to reconstruct a 3D scene from a collection of localised images, without any assumption on the nature of the scene nor viewpoints.
Our scene representation relies on the observation that grids of the I-NGP architecture define a unique mapping function $f:{\mathbb{R}}^{3}\rightarrow{\mathbb{R}}^{d}$ between the 3D scene and the latent space of dimension $d$ of the decoder. We then suppose that representing a scene at multiple scales is equivalent to defining the mapping function with different levels of detail (LOD). Unlike I-NGP which uses a bounded hierarchy of grids to define a mapping function with a non-uniform distribution of its degrees of freedom with respect to the latent space dimensions [^2] I-NGP latent space is defined as a concatenation of sub-spaces, each subspace being associated with a grid of a different resolution. The complexity of the resulting mapping function is thus non-uniform depending on the dimension of the latent space., we use an unbounded hierarchical grid pyramid to define recursively a mapping function with finer LOD (section 3.2).
Based on this multi-scale scene representation, we adapt the usual volumetric rendering by adjusting the LOD of the mapping function with respect to the pixel size and the distance of the sample to the camera (section 3.3).
To further improve the robustness of our reconstruction process, we introduce a continuous coarse-to-fine estimation of the mapping function LOD (section 3.4).
3.2 Scene Representation Architecture
As stated in section 2, we build our architecture with two objectives in mind: the natural production of continuous LOD, as well as a position and LOD-invariant decoder for the extensibility of our reconstruction.
Architecture-induced Level-Of-Detail. The first objective is to propose an architecture that introduces a continuous and non-destructive level of detail representation by design. As stated in section 3.1, we suppose that representing LOD in grid-based Neural Fields is equivalent to representing the mapping function between the scene space and the latent space with different LOD.
Inspired by Free Form Deformation, we propose to use an interpolative mapping function based on a lattice of control points. Under the existence assumption of a subdivision scheme $S$ that increases the lattice resolution while keeping the mapping function unchanged, our interpolative mapping function is defined recursively from the initial lattice $G_{0}$ (that defines explicitly the position of the control points).
$G_{N+1}\leftarrow S(G_{N})+\delta_{N+1}$ (1)
where $\delta_{N+1}$ is the deviation of the control points at level $N+1$ with respect to the control points defined by $S(G_{N})$.
Such representation provides several advantages. First, since the mapping function is constructed in a top-down manner, refining recursively one level of detail from the previous one, the hierarchy size is unbounded and adding a level of detail keeps the previous ones valid. Secondly, a continuous LOD representation can be easily obtained by incorporating a weighting factor $\alpha$ in the recursive process defined in eq. 1 is as follow:
$G_{L}\leftarrow S(G_{N})+\alpha*\delta_{N+1}$ (2)
with $L\in]N,N+1]$ and $\alpha=L-N$. Finally, as long as the subdivision scheme $S$ is derivable, the optimization of a fine LOD will optimize the coarser LOD since the gradient will propagate through the subdivision function.
@@ 3.2 Scene Representation Architecture
While different implementation of such mapping function is possible, the simplest one consists of using a regular 3D grid as control points lattice and linear interpolation among the cells of the grid. Since the subdivision function $S$ is also a linear interpolation, the latent vector associated with a point P in the scene for a level of detail $N$ is simply obtained by interpolating linearly the control points of the cell the point lies in for each level inferior or equal to $N$, and summing the results. Since linear interpolation is derivable, our mapping function is also optimizable. The summing of the interpolation is then comparable to a residual connection, as illustrated in figure 2 and giving the name of the architecture RING-NeRF for Residual Implicit Neural Grids. Those connections ensure the gradient flow over the hierarchy, making the reconstruction process more robust as will be demonstrated experimentally.
Invariance to position and LOD of Decoder. The second objective is to keep the latent space invariant to any information on position and LOD. Regarding the invariance to position, it implies that the decoder does not input any positional encoding [17], unlike [16]. To guarantee invariance to LOD, our architecture necessarily relies on a unique decoder (unlike [16]) that is fed a latent vector of any resolution which is not affected by any LOD downweighting (unlike [1, 2, 3]) nor augmented for instance by scale featurization (unlike [3]). Relating to the decoder architecture itself, the latent vector that outputs the grid is first normalized to improve convergence stability. Because a summation replaces the concatenation usually seen in grid-based architectures, the feature is shorter and could face problems for expressing high frequencies. It is thus projected into a higher dimension space with a Random Fourier Feature mapping [17] followed by a shallow MLP whose target is to output density after one layer and radiance at the end of this MLP based on the direction of observation.

### v3

# RING-NeRF : Rethinking Inductive Biases for Versatile and Efficient Neural Fields

@@ Abstract
Recent advances in Neural Fields mostly rely on developing task-specific supervision which often complicates the models. Rather than developing hard-to-combine and specific modules, another approach generally overlooked is to directly inject generic priors on the scene representation (also called inductive biases) into the NeRF architecture. Based on this idea, we propose the RING-NeRF architecture which includes two inductive biases : a continuous multi-scale representation of the scene and an invariance of the decoder’s latent space over spatial and scale domains. We also design a single reconstruction process that takes advantage of those inductive biases and experimentally demonstrates on-par performances in terms of quality with dedicated architecture on multiple tasks (anti-aliasing, few view reconstruction, SDF reconstruction without scene-specific initialization) while being more efficient. Moreover, RING-NeRF has the distinctive ability to dynamically increase the resolution of the model, opening the way to adaptive reconstruction. Project page can be found at : https://cea-list.github.io/RING-NeRF
Keywords:
NeRF 3D Reconstruction Implicit Neural Representation
Figure 1: RING-NeRF is a simple and versatile architecture which tackles many NeRF common issues such as robustness to distance of observation, few view supervision and lack of scene-specific initialization for SDF-based reconstruction. It provides on-par performances in terms of quality with SotA dedicated solutions [4, 25, 11] and and in terms of efficiency with fast methods [20, 26].
@@ 1 Introduction
Neural Radiance Fields (NeRF) have emerged as a novel method for representing 3D scenes using neural networks. In its original design [13], a simple multi-layer perceptron (MLP) is trained to reproduce a continuous 5D function that outputs the density and radiance emitted in each direction ($\theta$, $\phi$) at each point (x, y, z) in space. This approach inspired many works due to the impressive quality of its novel view synthesis as well as its simplicity. However, the so-called "vanilla" NeRF architectures converge very slowly as they use large deep neural networks. Instant-NGP [14] introduced a new architecture based on a hierarchical hash-grid pyramid of learnable feature codes describing the 3D scene and combined with a much shallower MLP, called decoder, that transforms the concatenation of the codes, interpolated from the grids, into density and radiance. Resulting in local updates, this method reduced the training duration from hours to minutes.
@@
Nature of scene - by transitioning from object-centric scenes to open unbounded scenes, using mostly space contraction [3];
Robustness - by managing free motion trajectories with variations of the observation distance while avoiding aliasing artefacts, mostly through the integration of Level of Detail (LOD) in the model [2, 4]; or by reducing drastically the number of supervised views through different kinds of regularization [25, 8, 15];
Extensibility - by shifting from a holistic and fixed reconstruction process to an incremental (extensibility in the number of views) and adaptive (extensibility in resolution) process, mainly through the use of a frozen decoder [29, 12].
However, most solutions solely focus on one of these limitations and introduce specific and complex mechanisms that both increase the computational cost and lessen the possibilities of combination. The ability to solve jointly these main issues is however essential in real-world applications which often require a robust and extensible reconstruction in an unknown and unbounded environment.
In this article, rather than introducing yet another heavy and task-specific solution, we propose RING-NeRF, a versatile and simple NeRF architecture by rethinking usual grid-based models to introduce two inductive biases. We first represent a 3D scene as a continuous multi-scale representation and also make the decoder’s latent space invariant in position and scale. Together, these two priors enable the production of intrinsic continuous LOD of the scene without explicit supervision. We demonstrate experimentally that, when combined with adapted cone casting and coarse-to-fine optimization, the resulting architecture is able to compete on several tasks with on-par quality performances with dedicated state-of-the-art solutions while improving speed, robustness and extensibility. The overall process is also simple, easy to implement and generic enough to be coupled with specific solutions. Our contributions can be summarized as follows:
@@
    an architecture that, by construction, represents the scene with a continuous and unbounded level of detail without the need for LOD-specific supervision and which permits resolution-adaptive reconstruction;
@@
    a distance-aware forward mapping compatible with scene contraction, that takes benefits of the continuous multi-scale representation with an adapted cone casting process to avoid aliasing artefacts when varying the observation distance.
@@
    a continuous coarse-to-fine reconstruction process that improves the convergence and stability (especially in challenging setups such as supervision with few viewpoints or no scene-specific initialization for SDF reconstruction).
2 Related Work
@@
From its original iteration [13], a majority of current research focuses on overtaking limitations of NeRF-based reconstruction in terms of adaptability to various natures of scene, robustness (to varying observation distances or limited amount of viewpoints) and reconstruction extensibility.
@@
Adaptability to various natures of the scene. The ability of Neural Fields to reconstruct various natures of scene depends on three factors. The first one is related to its architecture itself. Tri-plane architectures [7, 30, 5] are mostly designed for object-centric reconstruction (as they provide a higher density of information in the center of the scene) while vanilla and 3D grid-based NeRF are able to cover a wider variety of scenes, though they initially were still constrained to a limited volume. A second factor is related to the representation of the 3D space of the scene, especially to represent distant elements in open scenes. Some approaches use two different NeRF models to reconstruct separately the foreground and the background [28], whereas others apply space contraction to the 3D scene coordinates [28, 3, 4] to map the infinite scene volume into a bounded one. The last factor is related to the initialization of the Neural Fields. While the random initialization of density-based NeRF can adjust to almost any nature of scenes, the convergence of SDF-based (Signed Distance Function) Neural Fields is extremely sensitive to their parameters’ initialization, as stated in [1]. Current solutions rely on a scene-specific initialization (using an SDF field representing a sphere for outdoor scenes or an inverted sphere for indoor scenes), making them unable to adapt automatically to any scene.
@@
Robustness to observation distance variations. The initial NeRF model, as well as most of the subsequent works, relies implicitly on the hypothesis of a constant distance of observation to the scene. Indeed, the NeRF model provides a per-3D-point scene density representation and a rendering process which does not take into account the increasing volume covered by a pixel with respect to the distance to the cameras. As underlined in Mip-NeRF [2], this discrepancy induces artefacts such as over-contrasted images or aliasing phenomenon when the distance of observation differs from the ones used at reconstruction time. To avoid these artefacts, the rendering process needs to assess the density and color for a volume instead of a point. In the current state-of-the-art, two main approaches are used (or their combination, as in Zip-NeRF [4]). The first solution consists of representing the scene with different levels of detail (LOD), in order to vary the precision of the reconstruction based on the observation distance. This is usually done by using a LOD-aware latent space [6, 19, 21, 2, 4, 24], meaning that the LOD information is already encoded in the inputs of the MLP. This can be achieved by using a per-LOD decoder as in PyNeRF [21] or by incorporating LOD information in the latent feature as done in Zip-NeRF [4] and VR-NeRF [24]. One main flaw of these solutions is that they require the supervision of every used LOD which makes it impossible to vary the observation distances between the train views and novel synthesised views. A second approach consists of defining the latent representation of a volume as the mean of the latent features of the points included in the volume. It requires integrating the features over the volume, which can be achieved through convolutions for tri-plan representation [7, 30], or through super-sampling of the latent space for 3D grids as also done in Zip-NeRF [4]. However, these approaches lengthen the training and rendering processes as they increase the number of computations.
@@
Robustness to limited amount of viewpoints. By construction, NeRF is subject to the shape-radiance ambiguity [28]. If not enough supervision viewpoints are available, the optimization might overfit them while not providing consistent 3D reconstruction nor generalization to non-supervised viewpoints. A first family of solutions to overcome this limitation consists of regularizing the reconstruction process through additional losses, whether via geometric [15, 27, 23] or semantic [8] regularization. The second family of solutions relies on a progressive reconstruction of the details of the scene, as introduced in FreeNeRF [25] and Nerfies [17]. Restraining the ability of the model to reconstruct a complex scene at early stages enforces the consistency of the reconstruction over the different supervision viewpoints. This presents the advantage of bringing stability while keeping a fairly simple training process. However, these latter solutions mostly rely on Vanilla models for stability purposes and require long training duration.
@@
Reconstruction extensibility. Two kinds of extensibility should be distinguished: extensibility with respect to the number of views or the resolution. The first one is related to reconstructing new scene areas while keeping the previously reconstructed ones unchanged. The usual solution consists of using grid-based approaches with a pre-trained position-invariant decoder that is frozen during the reconstruction[29]. On the other side, the extensibility of the resolution consists in dynamically increasing the level of detail of a previously reconstructed scene. This problem is intractable for classic architectures such as vanilla NeRF and I-NGP since the number of layers (resp. grids) of those approaches cannot increase during the reconstruction. Some rare solutions, such as Neural Sparse Voxel Fields [12], combine a pre-trained decoder with a data structure allowing to dynamically allocate voxels to increase the reconstruction resolution. This is however a crucial stake, as being able to adapt the precision of the representation based on the complexity of the scene permits to optimize the computational efficiency (both in speed and memory requirements) with minimal loss in quality.
@@ 3 RING-NeRF
Rather than focusing on solving one specific problem of NeRF using complex mechanisms, we propose a simple architecture called RING-NeRF constructed with novel inductive biases to tackle NeRF’s common issues.
@@ 3.1 Overview
Figure 2: Overview of RING-NeRF: to render a pixel, the casted cone is sampled with cubes. Depending on the cube volume, the corresponding LOD of the scene is selected and the latent feature is computed using a weighted sum of the grid hierarchy. The density (or SDF) and color of the cube are first decoded from the latent feature with a tiny MLP and then integrated with other samples through volume rendering.
RING-NeRF relies on the classic NeRF [13] inverse rendering pipeline which is used to reconstruct a 3D scene from a set of localized frames. For a given image pixel (with its camera’s pose), a 3D ray is cast and the 3D scene representation is sampled at $N$ various locations along the ray. The resulting density $\sigma_{i}$ (or SDF converted to density [22]) and color $c_{i}$ of the samples are then combined with usual volume rendering techniques: $\hat{C}(r)=\sum_{i=0}^{N-1}T_{i}(1-exp(-\sigma_{i}\delta_{i}))c_{i}$ with $T_{i}=exp(\sum_{j=0}^{i-1}\sigma_{j}\delta_{j})$ and $\delta_{i}$ is the distance between samples. The parameters of the scene representation are then optimized by minimizing the MSE loss $||\hat{C}(r)-C(r)||^{2}$ between the rendered ray color and the ground truth pixel value, the rendering being differentiable [^1] With SDF, an additional Eikonal loss is also being optimized for SDF regularization..
The originality of RING-NeRF relies on its neural architecture, illustrated in figure 2, which is designed to represent the scene with a continuous and unbounded level of detail without the need for LOD-specific supervision (see section 3.2). We then use this LOD inductive bias to adjust the LOD of the samples with respect to the distance to the camera in contracted space for more accurate renderings (see section 3.3). Finally, combined with a continuous coarse-to-fine optimization, RING-NeRF results in a more robust reconstruction process with an intrinsic LOD extensibility property (see section 3.4).
3.2 NeRF Representation with Inductive Biases
Classic grid-based representations, such as the one introduced by I-NGP[14], rely on two key elements: a unique MLP decoder which transforms a latent space into an output space (color, density or SDF), and a 3D (hierarchical) grid of latent features that implicitly defines a unique mapping function from the scene space onto the decoder latent space. We can distinguish two different approaches used to extend the mono-scale representation to a multi-scale representation: conditioning with scale information the decoding process [4, 21] and/or defining separate per-LOD mapping functions [21]. In both cases, the training becomes more complex, leading to potential convergence issues as it relies solely on additional specific supervision. Instead, we propose to condition the mapping function itself with scale information and introduce inductive biases to guide the convergence.
Scene representation as multi-scale mapping function. As illustrated in Figure 2, the mapping function at level $N$ is controlled by a grid $G_{N}$ that is implicitly defined through a recursive refinement process over scale:
$\begin{cases}G_{0}\leftarrow\delta_{0}\\ G_{N}\leftarrow S(G_{N-1})+\delta_{N}\end{cases}$ (1)
@@ 3.2 NeRF Representation with Inductive Biases
where $\delta_{i}$ is the deviation defined by the i-th latent feature grid and S is the subdivision scheme consisting in increasing the resolution of $G_{N-1}$ up to the $N$-th grid resolution (with tri-linear interpolation). In practice, the latent vector associated with a point in the scene for a level of detail $N$ is simply obtained by interpolating linearly the point in the cell for each level inferior or equal to $N$, and summing the results. Since linear interpolation is differentiable, our mapping function is also optimizable. The summing of the interpolation is comparable to a residual connection and gives the name of the architecture RING-NeRF for Residual Implicit Neural Grids.
Such representation provides several advantages. First, since the mapping function is constructed in a top-down manner, refining recursively one level of detail from the previous one, its number of LOD is unbounded and adding a finer level of detail keeps the coarser ones valid. Secondly, because each $\delta_{i}$ only represents a deviation of the coarser mapping value $G_{i-1}$, a continuous LOD representation can be easily obtained by incorporating a weighting factor $\alpha\in[0,1]$ in the recursive process:
$G_{L}\leftarrow S(G_{N-1})+\alpha\delta_{N}$ (2)
with $L\in[N-1,N]$ and $\alpha=L-(N-1)$.
Spatial and scale invariance of decoder. Since our architecture does not rely on a decoder conditioned on position (unlike [19]) nor scale (unlike [2, 3, 4, 21]), the decoder latent space is invariant to translations and scale changes in the scene coordinate. This property makes RING-NeRF more suited for incremental reconstruction in both spatial and scale space, since it ensures that local updates in the spatial and scale domain of the scene can be achieved through the hierarchical grid. The decoder architecture is illustrated in figure 2 and further details can be found in the supplementary materials.
LOD inductive bias. During the reconstruction process, the gradients of the ray samples are backpropagated through the residual connections and aggregated for each grid level. Due to the pyramidal resolution of the hierarchical grid, a grid code at coarse levels influences a large scene volume and is thus supervised by more ray samples. Therefore, the gradient of a grid code increases as the level’s resolution decreases, and as long as the associated samples’ gradients are uniform. However, once the backpropagation through the residual connections reach a level whose associated samples have divergent enough gradients (meaning the error is finer than this level’s resolution), the result of the aggregation will be mitigated. Hence, the level with the maximum correction is always the coarser level where the samples’ gradients are still uniform. This property naturally induces corrections at the proper grid level and LOD. In figure 3, we illustrate this inductive bias by displaying different continuous LOD of a scene while the reconstruction is only supervised for its finest LOD. Not only this is a useful property when an unsupervised multi-scale representation is needed (as demonstrated in section 4.2), this also guarantees a more robust convergence (as illustrated in section 4.3 and section 4.4).
Figure 3: We demonstrate the LOD inductive bias by training our model with a hierarchy of $N=7$ levels where only the last level of the mapping function $G_{N}$ is supervised. We then compute renders at different levels $L\leq N$. Other examples (including of entire scenes) can be found in the supplementary materials.